# Not All Preference Pairs Are Created Equal: A Recipe for Annotation-Efficient Iterative Preference Learning

### v1

@@ Abstract
A Recipe for Annotation-Efficient Iterative Preference Learning
@@ 2 Preliminaries
In this section, we give a brief review of direct preference learning methods (§$\S$2.1) and iterative DPO (§$\S$2.2). Then we present the preliminary setup of this work (§$\S$2.3).
@@ 2.2 Online Iterative DPO
2.3 Experimental Setup
2.3.1 Synthetic Oracle
2.3.2 Training
2.3.3 Evaluation
@@ 3 Margin-based Selection within One Iteration
Instead of staring at a single margin value that provides little insight for data selection, we consider ranking a set of reward margins. We assume the highest- or lowest-ranking subset might be of use to find the worth-annotating instances. We do not have a formal theory to support this assumption, but there are some intuitions from two points of view, i.e., discriminative and generative, associated with the mixed nature of DPO that optimizes a pairwise discriminative function that is built upon the output logits of generative LLMs. We will show empirical evidence to support these intuitions in §$\S$3.4.
@@ 3.1 Strategy Variants
3.2 Experiment Setup
@@ 4 Training for Multiple Iterations
Given our findings in the single-iteration case in §$\S$3, we would like to know how well the winning strategies generalize to multiple iterations (Q4.1). Besides, recent work reported continuous improvements with multiple iterations for iterative DPO (Wu et al., 2024; Dong et al., 2024). They all adopted an evenly distributed strategy to allocate annotation budgets across iterations. A natural question is, whether the model benefits from training on more instances in earlier or later iterations, rather than always training on the same amount (Q4.2).
4.1 Experiment Setup
@@ 5.1 Iterative Preference Learning
Online iterative preference learning refers to the framework in which response pairs are sampled from the policy models and are then annotated to become the training data to continuously improve the policy model. Intuitively, it could mitigate the reward hacking issue or the distribution shift issue (Gao et al., 2022; Rafailov et al., 2024) Online iterative preference learning has been verified to be effective for alignment methods with explicit rewards (Bai et al., 2022; Touvron et al., 2023b) and for direct preference learning methods (Xu et al., 2024; Yuan et al., 2024; Xiong et al., 2024; Rosset et al., 2024; Wu et al., 2024; Swamy et al., 2024; Tran et al., 2023; Ye et al., 2024; Guo et al., 2024; Tajwar et al., 2024; Calandriello et al., 2024). Specifically, online direct preference learning was first presented in Xu et al. (2024). Dong et al. (2024) shows a systematic training pipeline and releases a strong policy checkpoint based on LLaMA-3-8B-base. Another line of work has investigated Nash equilibrium for LLM alignment, which is shown to natively support online iterative training by theory (Wu et al., 2024; Rosset et al., 2024; Munos et al., 2024).

### v2

@@ Abstract
A Recipe for Annotation-Efficient Iterative Preference Learning thanks: The work described in this paper is substantially supported by a grant from the Research Grant Council of the Hong Kong Special Administrative Region, China (Project Code: 14200719).
@@ 2 Preliminaries
In this section, we give a brief review of direct preference learning methods (§$\S$2.1) and iterative DPO (§$\S$2.2). Then we present the preliminary setup of this work (§$\S$3.2).
@@ 3 Margin-based Selection within One Iteration
Instead of staring at a single margin value that provides little insight for data selection, we consider ranking a set of reward margins. We assume the highest- or lowest-ranking subset might be of use to find the worth-annotating instances. We do not have a formal theory to support this assumption, but there are some intuitions from two points of view, i.e., discriminative and generative, associated with the mixed nature of DPO that optimizes a pairwise discriminative function that is built upon the output logits of generative LLMs. We will show empirical evidence to support these intuitions in §$\S$3.5.
@@ 3.1 Strategy Variants
3.2 Experimental Setup
3.2.1 Synthetic Oracle
3.2.2 Training
3.2.3 Evaluation
3.3 Empirical Workflow
@@ 4 Training for Multiple Iterations
Given our findings in the single-iteration case in §$\S$3, we would like to know how well the winning strategies generalize to multiple iterations (Q4.1). Besides, recent work reported continuous improvements with multiple iterations for iterative DPO (Wu et al., 2024; Dong et al., 2024). They all adopted an evenly distributed strategy to allocate annotation budgets across iterations. A natural question is, whether the model benefits from training on more instances in earlier or later iterations, rather than always training on the same amount (Q4.2). The background experimental setup follows that of §$\S$3.2.
4.1 Empirical Workflow
@@ 5.1 Iterative Preference Learning
Online iterative preference learning refers to the framework in which response pairs are sampled from the policy models and are then annotated to become the training data to continuously improve the policy model. Intuitively, it could mitigate the reward hacking issue or the distribution shift issue (Gao et al., 2022; Rafailov et al., 2024) Online iterative preference learning has been verified to be effective for alignment methods with explicit rewards (Bai et al., 2022; Touvron et al., 2023b) and for direct preference learning methods (Xu et al., 2024; Yuan et al., 2024; Xiong et al., 2024; Rosset et al., 2024; Wu et al., 2024; Swamy et al., 2024; Tran et al., 2023; Ye et al., 2024; Guo et al., 2024; Tajwar et al., 2024; Calandriello et al., 2024; Chen et al., 2024). Specifically, online direct preference learning was first presented in Xu et al. (2024). Dong et al. (2024) shows a systematic training pipeline and releases a strong policy checkpoint based on LLaMA-3-8B-base. Another line of work has investigated Nash equilibrium for LLM alignment, which is shown to natively support online iterative training by theory (Wu et al., 2024; Rosset et al., 2024; Munos et al., 2024).